OpenStudy (ammarah):

Solve the system using any algebraic method!!!!!! HELP I BEG U 2x+y-z=-1 x+2y+2z=10 2x+z=3

OpenStudy (ammarah):

@Zale101

OpenStudy (zale101):

first, pick an equation

OpenStudy (zale101):

then, solve for any letter that u choose

OpenStudy (zale101):

tell me the equation and the letter u choose. then we can solve it together :)

OpenStudy (zale101):

@Ammarah, enti ma3y?

OpenStudy (ammarah):

ok the first and second one

OpenStudy (ammarah):

ughh i dont get itt.....

OpenStudy (ammarah):

can u please show me all the stepss...

OpenStudy (ammarah):

yess i know....but i dont know what to do after that...

OpenStudy (ammarah):

thats the first equation though........

OpenStudy (ammarah):

please solve it for me i dont get it....after i look over what u did ill understandd.....

OpenStudy (zale101):

2x+y-z = -1 we are finding the variable for y. Move the 2x to the right hand side by subtracting 2x from both sides. the equation reads: y = -1-2x+z when u isolate the y, you have to divide both sides of the equation by the other variables

OpenStudy (zale101):

the y has to be on the left side of the equation since we're solving for y

OpenStudy (ammarah):

now what?

OpenStudy (zale101):

do the next equation x+2y+2z=10 you have to substitute the expression you get for that in each of the other equations and therefore u simplify

OpenStudy (ammarah):

can u like tell me the answer...then ill try to solve it and plug it in...

OpenStudy (ammarah):

like can uy draw out the steps?

OpenStudy (ammarah):

yeah sure ill wait thanks...

OpenStudy (ammarah):

@zale r u still there?